# Exploring Parity Challenges in Reinforcement Learning through Curriculum Learning with Noisy Labels

### 08 Dec 2023 v1

@@ 1 Introduction
Prior work has concluded that learning a parity function with neural networks from long bitstrings sampled from a uniform distribution is a challenge Cornacchia and Mossel (2023); Abbe et al. (2023); Daniely and Malach (2020); Shalev-Shwartz et al. (2017). Recent studies have explored ways to build non-uniform distributions to make it learnable Cornacchia and Mossel (2023); Abbe et al. (2023); Daniely and Malach (2020). We will summarise their results in Section 2.
@@
The structure of this paper is as follows. Section 2 summarises recent studies on the intricacies of modelling a parity function using various neural networks. Section 3 explains our experiment setup and presents our results on the noisy label problem in a simulated reinforcement learning setting. This paper is concluded in Section 4.
@@ 2 Related Work
Learning the parity function using a neural network is a trivial task on short bitstrings. One of our experiments showed that a single-layer Long Short-Term Memory (LSTM) model with 16 nodes could model a perfect parity function on uniformly sampled bitstrings of length 20. A request of research  [^1] https://openai.com/research/requests-for-research-2 released by OpenAI proposed a question with regards to using LSTM to model the parity function under two different sets of bitstrings. One database consists of 100,000 random bitstrings of a fixed length of 50, whereas the other one contains the same number of bitstrings, but the length of each bitstring is independently and randomly chosen between 1 and 50. The results were with the first database, the LSTM model failed to learn the parity function, but it succeeded with the second dataset. Additionally, on the bitstrings with varying lengths, the LSTM model tends to converge faster if it was fit in shorter bitstrings than longer ones at the early training stage. This suggests that the performance of the neural network in modelling the parity function is influenced not only by the length of the bitstrings but also by the distribution from which they are sampled.
@@ 2.2 Learning parity function with non-uniform data
Although learning parity function under uniform data distribution on long bitstrings faces strict constraints, some prior studies have shown that when the bitstrings are not sampled uniformly, neural networks trained by gradient descent methods can model the parity function efficiently with the data sampled from certain distributions. Cornacchia and Mossel (2023) proposed a new way to build the dataset in which the bitstrings are sampled from two or more product distributions and the neural network is trained using curriculum learning (CL), leading to a significant reduction in the computational cost of learning the parity function. Abbe et al. (2023) further proved that learning the parity function in the curriculum learning fashion provides certain advantages over the standard training method where data are sampled uniformly. In CL, the model is presented with spare examples at the early stage of training after which the whole dataset is presented to the model. This enables the model to learn some knowledge in the beginning and primes it to learn from complex data.
@@
Pondernet is a new type of neural network tested against modelling the parity function Banino et al. (2021). To make the parity function learnable, in their setting, the bitstrings of length $n$ are sampled in a way that a random number of bits from 1 to $n$ were randomly set to 1 or -1 and the remaining bits were set to 0. This approach establishes a latent curriculum where the model is presented with a mix of spare and dense bitstrings in each training batch. Here, the density of a bitstring is measured by Hamming weight, where the hardness of a bitstring is defined by the number of 1 and -1 bits in it, analogous to the measure used in Abbe et al. (2023).
@@ 3.1 Learning parity function from latent curriculum
Our results align with the results from Daniely and Malach (2020) that showed on bitstrings sampled from normal distribution, neural networks on longer bitstrings require more training steps to converge. While the training process with the bitstrings constituting a latent curriculum differs from the CL where only spare samples are presented at the early stage of training as each batch of samples fitting into the NN consists of both spare and dense samples, our results on the latent curriculum can be extended to a CL setting that mirrors the way the data samples are presented to NNs in a self-play RL algorithm.

### 14 Jan 2024 v2

@@ 1 Introduction
Prior work has concluded that learning a parity function with neural networks from long bitstrings sampled from a uniform distribution is a challenge Cornacchia and Mossel (2023); Abbe et al. (2023); Daniely and Malach (2020); Shalev-Shwartz et al. (2017); Raz (2018). Recent studies have explored ways to build non-uniform distributions to make it learnable Cornacchia and Mossel (2023); Abbe et al. (2023); Daniely and Malach (2020). We will summarise their results in Section 2.
@@
The structure of this paper is as follows. Section 2 summarises recent studies on the intricacies of modelling a parity function using various neural networks. Section 3 explains our experiment setup and presents our results on the noisy label problem in a simulated reinforcement learning setting. This paper is concluded in Section 4. The codes and the data used in this paper are publicly available  [^1] GitHub. https://github.com/sagebei/Exploring-Parity-Challenges-in-Reinforcement-Learning.
@@ 2 Related Work
Learning the parity function using a neural network is a trivial task on short bitstrings. One of our experiments showed that a single-layer Long Short-Term Memory (LSTM) model with 16 nodes could model a perfect parity function on uniformly sampled bitstrings of length 20. A request of research  [^2] OpenAI. https://openai.com/research/requests-for-research-2 released by OpenAI proposed a question with regards to using LSTM to model the parity function under two different sets of bitstrings. One database consists of 100,000 random bitstrings of a fixed length of 50, whereas the other one contains the same number of bitstrings, but the length of each bitstring is independently and randomly chosen between 1 and 50. The results were with the first database, the LSTM model failed to learn the parity function, but it succeeded with the second dataset. Additionally, on the bitstrings with varying lengths, the LSTM model tends to converge faster if it was fit in shorter bitstrings than longer ones at the early training stage. This suggests that the performance of the neural network in modelling the parity function is influenced not only by the length of the bitstrings but also by the distribution from which they are sampled.
@@ 2.2 Learning parity function with non-uniform data
Although learning parity function under uniform data distribution on long bitstrings faces strict constraints, some prior studies have shown that when the bitstrings are not sampled uniformly, neural networks trained by gradient descent methods can model the parity function efficiently with the data sampled from certain distributions. Cornacchia and Mossel (2023) proposed a new way to build the dataset in which the bitstrings are sampled from two or more product distributions and the neural network is trained using curriculum learning (CL), leading to a significant reduction in the computational cost of learning the parity function. Abbe et al. (2023) further proved that learning the parity function in the curriculum learning fashion provides certain advantages over the standard training method where data are sampled uniformly. In CL, the model is presented with sparse examples at the early stage of training after which the whole dataset is presented to the model. This enables the model to acquire some knowledge in the beginning and primes it to learn from complex data.
@@
Pondernet is a new type of neural network tested against modelling the parity function Banino et al. (2021). To make the parity function learnable, in their setting, the bitstrings of length $n$ are sampled in a way that a random number of bits from 1 to $n$ were randomly set to 1 or -1 and the remaining bits were set to 0. This approach establishes a latent curriculum where the model is presented with a mix of sparse and dense bitstrings in each training batch. Here, the density of a bitstring is measured by Hamming weight, where the hardness of a bitstring is defined by the number of 1 and -1 bits in it, analogous to the measure used in Abbe et al. (2023).
@@ 3.1 Learning parity function from latent curriculum
Our results align with the results from Daniely and Malach (2020) that showed on bitstrings sampled from normal distribution, neural networks on longer bitstrings require more training steps to converge. While the training process with the bitstrings constituting a latent curriculum differs from the CL where only sparse samples are presented at the early stage of training as each batch of samples fitting into the NN consists of both sparse and dense samples, our results on the latent curriculum can be extended to a CL setting that mirrors the way the data samples are presented to NNs in a self-play RL algorithm.